A coil of Nichrome wire is 29.0 m long. The wire has a diameter of 0.460 mm and is at 20.0°C. 


(a) If it carries a current of 0.460 A, what is the magnitude of the electric field in the wire? 

(b) If it carries a current of 0.460 A, what is the power delivered to it? 

(c) If the temperature is increased to 360ºC and the potential difference across the wire remains constant, what is the power delivered?
Answer #1

Resistance, R = pL/A

R = (1.1 x 10^-6 x 29)/pi (0.23 x 10^-3)^2 = 191.95 ohm

Voltage, V = 0.46 x 191.95 = 88.3 V

A)

Electric field, E = V/d = 88.3/29 = 3.04 V/m

B)

Power, P = I^2R = 0.46^2 x 191.95 = 40.62 W

C)

R = 191.95 (1 - 0.4 x 10^-3(360 - 20)) = 165.84 ohm

Power, P = 0.46^2 x 165.84 = 35.09 W

Comment in case any doubt please rate my answer....